OpenStudy (anonymous):

Simplify: (5a^2/3)(4a^3/2)

OpenStudy (anonymous):

No, they're multiplied not divided. Other than that, yes

OpenStudy (saifoo.khan):

RIght,

OpenStudy (saifoo.khan):

\[\Large 20a^\frac{13}{6}\]

jimthompson5910 (jim_thompson5910):

\[\Large \left(5a^{\frac{2}{3}}\right)\left(4a^{\frac{3}{2}}\right)\] \[\Large \left(5 \times 4\right)\left(a^{\frac{2}{3}}\times a^{\frac{3}{2}}\right)\] \[\Large 20a^{\left(\frac{2}{3}+\frac{3}{2}\right)}\] \[\Large 20a^{\frac{13}{6}}\] So \[\Large \left(5a^{\frac{2}{3}}\right)\left(4a^{\frac{3}{2}}\right)=20a^{\frac{13}{6}}\]

OpenStudy (anonymous):

Thank you for the steps, that helps a lot!